OpenStudy (anonymous):

Each year about 9*10^7 babies are born in the world. A year contains 5*10^5 minutes. How many babies are born each minute?

OpenStudy (anonymous):

no of babies born per minute = (9*10^7)/(5*10^5) = 9/5* 10^(7-5) = 1.8*102

OpenStudy (anonymous):

That's what I got, but it's not one o the choices....the choices are A. 530 B. 1.7*10^2 C. 17*10^2 D. 1.7*10^12